Owl shift, 16 Sep 1999 (fwd)

Hogan Nguyen (hogann@fnal.gov)
Thu, 16 Sep 1999 06:41:40 -0500 (CDT)

Shift Summary for OWL shift, 16/SEP/99

On Shift: Jianbo, Hogan

Hours of beam delivered at all intensities: 0.0
Hours of beam delivered at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 7.0

Accounting of lost time:

========================================================================
Runs ended this shift: None

Tapes brought to FCC this shift: None

========================================================================

Accesses during shift: Open but supervised access

========================================================================
Overview of shift: DC work.

========================================================================
Description of Shift:

All He Bags reinflated.
Helium tanks 17+18 are now in use.
DC3 noise debugging finished.
All DC4 DCORs plugged in. Stable so far.

DC HVs 100 volts below nominal for now.
DC4 HV turned off in anticipation of DC4 realignment later this
morning.

========================================================================
Follow up on Notes from Previous Shift

========================================================================
Items to Follow up on Next Shift

Continue to keep an eye on KTeV hall temp; Et is still off in anticipation
of chilled water power down early tommorrow AM.

Continue surveys in morning, perhaps salvage GC of DC gas(?), reduce
aperture of 90 degree monitor. Debug DC4 DCORs latches
for DC3 & DC4. More fruit debugging, testing of E799 readout. After
survey, reassemble NM2 vacuum line, bring up TRD HV, move mask, REG out.